Moore Company has the following partial list of account balances at year-end:

Account payable $1,500

Accounts receivable 4,600

Cost of good sold 3,200

Cash 23,000

Taxes payable 10,000

Land 25,000

Notes payable( Due in6 months) 1.000

Salaries payable 900

Inventory 4,300

Requirements :

A.compute the quick ratio

B.Determine the amount of work capital

C. Assuming that cash is used pay the balance due on accounts payable

1.new compute the quick ratio

2.compute the new amount of work capital

D. Compute the accounts payable turnover ratio(use year end amounts)
Answer #1

Answer A:

Quick assets = Cash and cash equivalents + Marketable securities + Accounts receivable

= 23000 + 0 + 4600

=$27,600

Current liabilities = Account payable + Taxes payable + Notes payable (Short - term) + Salaries payable

= 1500 + 10000 + 1000 + 900

= $13,400

Quick ratio = Quick assets / Current liabilities = 27600 / 13400 = 2.06

Quick ratio = 2.06

Answer B:

Current assets = 4600 + 23000 + 4300 = $31,900

Working capital = Current assets - Current liabilities = 31900 - 13400 = $18,500

Working capital = $18,500

Answer C (1):

Assuming that cash is used pay the balance due on accounts payable:

Cash balance = 23000 - 1500 = $21,500

New Quick assets = 21500 + 4600 = $26,100

New current liabilities =10000 + 1000 + 900 = $11,900

New quick ratio = 26100 / 11900 = 2.19

New quick ratio = 2.19

Answer C (2):

New current assets = 4600 + 21500 + 4300 = $30,400

New amount of work capital = 30400 - 11900 = $18,500

New amount of Working capital = $18,500

Answer C (3):

Accounts payable Turnover ratio = Cost of goods sold / Year end accounts payable = 3200 / 1500 = 2.13333

Accounts payable Turnover ratio = 2.13